# The multimessenger view of Pulsar Timing Array black holes with the Horizon-AGN simulation

### v1

@@
22institutetext: Institut d’Astrophysique de Paris, UMR 7095, CNRS and Sorbonne Université, 98 bis boulevard Arago, 75014 Paris, France 33institutetext: IRAP, Université de Toulouse, CNRS, UPS, CNES, 9 Avenue du Colonel Roche, BP 44346, 31028 Toulouse Cedex 4, France 44institutetext: LERMA, Sorbonne Université, Observatoire de Paris, PSL research university, CNRS, 75014 Paris, France 55institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Edinburgh EH9 3HJ, UK
The multimessenger view of Pulsar Timing Array black holes with the Horizon-AGN simulation
@@
We use the Horizon-AGN cosmological simulation to study the properties of supermassive black hole binaries (MBHBs) contributing most to the gravitational wave background (GWB) signal expected in the pulsar timing array (PTA) band. We develop a pipeline to generate realistic populations of MBHBs, allowing us to estimate both the characteristic strain and GWB time series observable by PTA experiments. We identify potential continuous wave (CW) candidates standing above the background noise, using toy PTA sensitivities representing the current EPTA and future SKA. We estimate the probability of detecting at least one CW with signal-to-noise ratio >3>3 to be $4\%$ ($20\%$) for EPTA (SKA)-like sensitivities, assuming a 10-year baseline. We find the GWB to be dominated by hundreds to thousands of binaries at redshifts in the range 0.0510.05-1, with chirp masses of $10^{8.5}-10^{9.5}\,\,{\rm M_{\odot}}$, hosted mainly in quiescent massive galaxies residing in halos of mass $\sim 10^{13}\,\,{\rm M_{\odot}}$. CW candidates have larger masses, lower redshifts and are found in even more massive halos, typical of galaxy groups and clusters. The majority of these systems would appear as AGN rather than quasars, because of their low Eddington ratios. Nevertheless, CW candidates with $f_{\rm Edd}>10^{-3}$ can still outshine their hosts, particularly in radio and X-ray bands, suggesting them as the most promising route for identification. Our findings imply that optical and near-infrared searches based on light curve variability are challenging and biased toward more luminous systems. Finally, we highlight important caveats in the common method used to compare PTA observations with theoretical models. We find that GWB spectral inferences used by PTAs could be biased toward shallower slopes and higher amplitudes at $f=1/\rm yr$, thereby reducing the apparent tension between astrophysical expectations and PTA observations.
@@ 1 Introduction
Recently, various PTA experiments have presented first observational evidence for a GW signal (NANOGrav, Agazie et al. (2023a); EPTA & InPTA, EPTA Collaboration et al. (2023b); PPTA, Reardon et al. (2023); CPTA Xu et al. (2023) and MeerKAT, Miles et al. (2024)). This signal shows both spatial and temporal correlations among pulsars, consistent with those expected from a stochastic GW background (Hellings & Downs, 1983). The low signal to noise ratio of the detection makes the interpretation difficult, but one possibility is the non-coherent superposition of GWs from a population of massive black hole binaries (MBHBs) with separations of the order of milliparsecs. Indeed, inference of the GW background spectrum using PTA data confirms that the observed signal is broadly consistent with expectations from the MBHB population (Agazie et al., 2023c; EPTA Collaboration et al., 2024a), although the data tend to favor a larger amplitude than theoretically predicted. A possibility is that the signal emanates from individually resolvable MBHBs which are particularly close and heavy, therefore standing above the background signal e.g. Mingarelli et al. (2017); Agazie et al. (2023b). However, so far, this scenario is not favored by current PTA datasets (EPTA Collaboration et al., 2024b; Agazie et al., 2023b). But un-modeled noise or too agnostic priors for the noise parameters could also explain part of the tension (Zic et al., 2022; Goncharov et al., 2024; van Haasteren, 2025). In this work, we show that a too direct comparison between the theoretical power spectrum predicted by cosmological simulations (Kelley et al., 2017a; Sykes et al., 2022; Chen et al., 2025) or semi-analytical models (Wyithe & Loeb, 2003; McWilliams et al., 2014; Bonetti et al., 2018) and the data-inferred power spectra can lead to misleading conclusions about potential tensions between the two. In particular, the impact of the finite duration of the datasets on spectral inference should be considered more carefully [^1] After the submission of this paper, Crisostomi et al. (2025) also investigated the impact of improperly accounting for data windowing in PTA spectral inference..
@@
In the absence of a stronger GW detection, complementary data are necessary to understand the origin of the signal. Cosmological simulations can be used to understand the prevalence and nature of MBHBs in the Universe and identify which of these would be detectable with the PTAs and either contribute to the GW background and/or be detectable as individual objects. Electromagnetic observations of these objects would help confirm the nature of the GW sources, but simulations can help us better understand where to search and provide the best strategy for the wavelength ranges to target (see e.g. Cella et al., 2025). With such prior knowledge, the multi-messenger detection of MBHBs using PTAs, expected to be challenging (Charisi et al., 2022; Petrov et al., 2024), can be facilitated (Liu & Vigeland, 2021; Truant et al., 2025).
@@
In this paper, we use the Horizon-AGN cosmological simulation to identify and study the MBHBs that contribute most to the GW signal in the PTA band, either by contributing to the background signal or as resolvable continuous wave sources. We study the properties of these MBHBs along with those of their host galaxies. We determine the expected electromagnetic emission from radio to X-ray wavelengths, which we validate by comparing with the properties of MBHB candidates, and identify the types of searches that would have the best chance of detecting the electromagnetic counterparts of the PTA sources. Finally, we also investigate the commonly used methodology for comparing PTA observations with the predicted characteristic strain signal, highlighting its potential limitations when used to interpret the origin of the observed signal.
@@
The paper is organized as follows. Section 2 presents the Horizon-AGN simulation used in our study. In Section 3, we present how we build our catalog of merging MBHBs from the simulation and model their sub-parsec dynamics in post-processing. In Section 4, we present how we generate realizations of the inspiralling MBHB population and infer the associated characteristic strain signal. We present the properties of the most contributing binaries and estimate their number in the PTA band. In Section 5, we compare the standard methodology used to convert the GW strain signal into timing residuals measured by PTAs, with a more accurate method, and highlight the potential limitations of the former. Using two toy PTA sensitivities, we also identify potentially resolvable individual binaries, for which we analyze their properties and those of their host galaxies in Section 6, along with their electromagnetic properties, and evaluate their potential detectability in Section 7. We summarize our findings in Section 8.
@@ 2 The Horizon-AGN simulation
The gas in Horizon-AGN has an equation of state for an ideal monoatomic gas with an adiabatic index of $\gamma_{\rm ad}=5/3$. Gas cooling is modeled down to a floor temperature of $10^{4}\,\rm K$ with curves from Sutherland & Dopita (1993). Heating of the gas from a uniform UV background starts after redshift $z_{\rm reion}=10$ following the prescription in Haardt & Madau (1996). In regions which exceed a gas hydrogen number density threshold of $n_{0}=0.1\,\rm H\,cm^{-3}$, the star formation is triggered in a Poisson random process (Rasera & Teyssier, 2006; Dubois & Teyssier, 2008) following the Schmidt relation with a constant star formation efficiency $\varepsilon_{*}=0.02$ (Kennicutt, 1998; Krumholz & Tan, 2007). Mechanical energy injection from Type Ia SNe, Type II SNe and stellar winds is included assuming a Salpeter (1955) initial mass function with cutoffs at $0.1\,{\rm M_{\odot}}$ and $100\,{\rm M_{\odot}}$.
@@
Massive black holes and their feedback follow the numerical implementation of Dubois et al. (2012). They are seeded in cells where the gas density is larger than $n_{0}$, and where the gas velocity dispersion is larger than $100\,\rm km\,s^{-1}$. The MBH seeding is stopped at $z=1.5$. To avoid formation of multiple MBHs in the same galaxy, an exclusion radius of 50 comoving kpc is imposed in the seeding process. To compensate the inability to capture the multiphase nature of the interstellar gas, the MBH accretion rate is set to be the Bondi-Hoyle-Littleton rate multiplied by a factor $\alpha_{\rm b}=(n/n_{0})^{2}$ when $n>n_{0}$ and $\alpha_{\rm b}=1$ otherwise (Booth & Schaye, 2009). The effective accretion rate onto MBHs is capped at the Eddington luminosity with a radiative efficiency of 0.1. 15% of the MBH emitted energy is isotropically coupled to the gas within $4\Delta x$ as thermal energy for luminosities above 1% of the Eddington luminosity. If the 1% of the Eddington luminosity threshold is not reached, then the feedback takes a mechanical form instead, with 100% of the power injected into a bipolar cylindrical jet with a radius of $\Delta x$ and a height of $2\Delta x$, at a velocity of $10^{4}\,\rm km\,s^{-1}$.
@@
Instead of constantly repositioning MBHs at the minimum of the local potential, which causes unnatural dynamics (Tremmel et al., 2015), gas dynamical friction is exerted on the MBH to avoid spurious motions due to finite force resolution effects (see Dubois et al., 2013 for additional details). The boost factor used here is the same as the boost factor $\alpha_{\rm b}$ for accretion. The equation of gas dynamical friction is $F_{\rm DF}=f_{\rm gas}4\pi\alpha_{\rm b}\rho_{\rm gas}(GM_{\rm BH}/\bar{c}_{s})^{2}$, where $G$ is the gravitational constant, $\rho_{\rm gas}$ is the mass-weighted mean gas density within a sphere of radius $4\,\Delta x$ and $f_{\rm gas}$ is a factor function of the mach number ${\mathcal{M}}=\bar{u}/\bar{c}_{s}$, which accounts for the extension and shape of the wake (Ostriker, 1999). $f_{\rm gas}$ is in a range between 0 and 2 for an assumed Coulomb logarithm of 3 (Chapon et al., 2013; Lescaudron et al., 2023).
@@
In Horizon-AGN the dark matter haloes and galaxies are identified with the AdaptaHOP halo finder (Aubert et al., 2004). The density field used in AdaptaHOP is smoothed over 20 particles. The identification density threshold is set to be 178 times the average total matter density. On top of the density criteria, either 50 dark matter particles or 50 star particles are required for a dark matter halo or galaxy identification. The center of haloes and galaxies are estimated using the shrinking sphere approach proposed by Power et al. (2003).
@@ 3 Selection and modeling of MBH binaries in Horizon-AGN
A catalog of MBH mergers is generated using the post-processed simulation data. To find merging MBHs, we search through the data of all MBHs at each coarse time step ($\sim 0.6-0.7$ Myr) of the Horizon-AGN simulation. Note that this time step is much shorter than the time step over which a full output is saved, which is $\sim 150$ Myr. Each MBH is assigned a unique ID number when seeded. This ID is carried by the MBH through the simulation until it numerically merges with another MBH at a separation of $4\Delta x$. The merged MBH inherits the ID of the primary black hole in the merger, while the secondary MBH ID is erased after the merger. By searching for disappearing MBH IDs, we can find the corresponding MBH mergers. Possible spurious numerical mergers are filtered out by selecting only MBH pairs within $\max(2R_{\rm eff},4\Delta x)$ from the galaxy centers, where $R_{\rm eff}$ is the effective radius of the galaxy measured as its projected 2-dimensional radius containing half of the galaxy stellar mass. This check is performed for the primary MBH at the outputs before the merger and for the merged MBH at the output after the merger: the reason is that the MBHs are likely to move relative to the galaxy center in the $150$ Myr between outputs.
@@
First, we added a dynamical friction phase from the position of the MBHs when they are numerically merged down to the point when they become gravitationally bound. The dynamical friction timescale is estimated assuming the MBH is in an isothermal sphere, considering only the stellar component of the galaxy and including a factor $0.3$ to account for typical orbits being non-circular:
@@
where $M_{\rm BH}$ is the black hole mass, $\sigma_{\star}$ is the central stellar velocity dispersion approximated as $(0.25GM_{\star}/R_{\rm eff})^{1/2}$, $\Lambda=\ln(1+M_{\star}/M_{\rm BH})$, with $M_{\star}$ being the total stellar mass of the galaxy hosting the MBH at the output before the numerical merger, and $d$ its distance from the galactic center. As confirmed by Li et al. (2020a), Li et al. (2020b), and Chen et al. (2022), the stellar dynamical friction dominates over the dynamical friction from gas or that from dark matter. We calculated the dynamical friction delay time of both MBHs in a pair and used the longer, which is normally associated with $M_{2}$.
We also accounted for the shrinking of the binary orbit until coalescence via stellar hardening, viscous torques in a circumbinary disc, and emission of gravitational waves, following Sesana & Khan (2015) and Dotti et al. (2015). The binary evolution timescale to coalescence was taken to be the minimum between the two following equations:
@@
Where $q=M_{2}/M_{1}$ is the mass ratio of the MBHB. In Equation 2, $\sigma_{\rm inf}$ and $\rho_{\rm inf}$ are the velocity dispersion and stellar density at the sphere of influence, defined as the sphere containing twice the binary mass in stars:
@@
where we have continued to assume a singular isothermal sphere power-law density profile with slope 2-2, and $a_{\rm gw}$ is the separation at which the binary spends most of the time (see Sesana & Khan, 2015):
@@
The spatial resolution of the Horizon-AGN simulation does not provide reliable information on the density profile at such small orbital separations. However, it was demonstrated in (Volonteri et al., 2020) that the choice of a single isothermal sphere is in very good agreement with the densities measured in observations of local galaxies. In Equation 3, $\varepsilon_{0.1}$ is the radiative efficiency normalized to 0.1. We followed Dotti et al. (2015) in selecting $a_{\rm i}=GM_{12}/2\sigma_{\star}^{2}$ and $a_{\rm c}=1.9\times 10^{-3}(M_{12}/10^{8}\,{\rm M_{\odot}})^{3/4}\,{\rm pc}$.
To model the eccentricity evolution in post-processing, we first assign an eccentricity to each binary when reaching the influence radius $r_{\rm inf}$ according to the eccentricity distribution in the right panel of Figure 6 in Li et al. (2020a). The eccentricity is then evolved under loss-cone scattering, viscous drag from the circumbinary disk, and GW emission, from the influence radius to the final coalescence. The orbit hardening and eccentricity evolution due to the loss-cone scattering is described by:
@@
where $f_{\rm orb}$ is the orbital frequency, and $H$ and $K$ are numerical factors from three-body scattering experiments (Quinlan, 1996; Sesana et al., 2006).
The viscous drag due to the circumbinary disk is included when the separation between the binary is below $1\,\rm pc$. Haiman et al. (2009) described how the binary orbit embedded in a circumbinary (Shakura & Sunyaev, 1973) $\alpha$-disk evolves due to viscous drag, and how this evolution depends on the different physical conditions within the disk.
According to Haiman et al. (2009), there are different regimes for a MBHB in a gap-opened, $\alpha$-disk depending on: (1) whether the radiation pressure or gas pressure balance the vertical gravity ($r^{\rm gas/rad}$); (2) whether the opacity is dominated by electron scattering or free-free absorption ($r^{\rm es/ff}$); (3) whether the binary is massive enough compared to the local disk mass ($M_{\rm 2}$-dominated or disk-dominated). The characteristic radii are defined as:
@@
According to Shapiro & Teukolsky (1983), a disk can be divided into three regions: (i) inner region ($r<r^{\rm gas/rad}$) with radiation dominated pressure and electron-scattering dominated opacity; (ii) middle region ($r^{\rm es/ff}>r>r^{\rm gas/rad}$) with gas dominated pressure and electron-scattering dominated opacity; (iii) outer region ($r>r^{\rm es/ff}$) with gas dominated pressure and free-free scattering dominated opacity. And within each of those three region, there are two possibilities: 1) $M_{\rm 2}$-dominated region ($r<r^{\rm\nu/s}$); 2) disk-dominated region ($r>r^{\rm\nu/s}$). The $r^{\rm\nu/s}$ in three regions of an $\alpha$-disk are defined in Haiman et al. (2009) as:
@@
Thus, there are six regimes and their orbital frequency evolution rate are listed below.
@@
if $r^{\rm gas/rad}<r<r^{\rm es/ff}$ and $r>r_{\rm mid}^{\rm\nu/s}$; where $r_{\rm 3}$ is the orbital semi-major axis in units of $10^{\rm 3}R_{\rm sch}$, $q_{\rm s}=4q/(1+q)^{\rm 2}$ is the symmetric mass ratio. Note that this prescription implies that the MBHB orbit always shrinks under the influence of viscous drag, especially in the presence of stellar hardening suggested by some most recent simulations (Cuadra et al., 2009; Roedig et al., 2012; Bortolas et al., 2021; Franchini et al., 2021; Amaro-Seoane et al., 2023)

### v2

# The multi-messenger view of pulsar timing array black holes with the Horizon-AGN simulation

@@
22institutetext: Institut d’Astrophysique de Paris, UMR 7095, CNRS and Sorbonne Université, 98 bis boulevard Arago, 75014 Paris, France 33institutetext: IRAP, Université de Toulouse, CNRS, UPS, CNES, 9 Avenue du Colonel Roche, BP 44346, 31028 Toulouse Cedex 4, France 44institutetext: Institute for Astronomy, University of Edinburgh, Royal Observatory, Edinburgh EH9 3HJ, UK
The multi-messenger view of pulsar timing array black holes with the Horizon-AGN simulation
@@
We used the Horizon-AGN cosmological simulation to study the properties of supermassive black hole binaries (MBHBs) with the largest contribution to the gravitational wave background (GWB) signal expected for the pulsar timing array (PTA) band. We developed a pipeline to generate realistic populations of MBHBs, which enabled us to estimate both the characteristic strain and GWB time series observable by PTA experiments. We identified potential continuous wave (CW) candidates standing above the background noise, using toy PTA sensitivities representing the current EPTA and future SKA. We estimated the probability of detecting at least one CW with a signal-to-noise ratio of S/N >3>3 to be $4\%$ ($20\%$) for EPTA (SKA)-like sensitivities, assuming a ten-year baseline. We found that the GWB is dominated by hundreds to thousands of binaries at redshifts in the range 0.0510.05-1, with chirp masses of $10^{8.5}-10^{9.5}\,\,{\rm M_{\odot}}$, primarily hosted in quiescent massive galaxies residing in halos of mass $\sim 10^{13}\,\,{\rm M_{\odot}}$. CW candidates have larger masses and lower redshifts, and they tend to be found in even more massive halos, typical of galaxy groups and clusters. The majority of these systems would appear as active galactic nuclei (AGNs), rather than quasars, because of their low Eddington ratios. Nevertheless, CW candidates with $f_{\rm Edd}>10^{-3}$ can still outshine their hosts, particularly in radio and X-ray bands, suggesting that they could serve as the most promising route for identification. Our findings imply that optical and near-infrared (NIR) searches based on light curve variability are challenging and biased toward more luminous systems. Finally, we highlight important caveats in the common method used to compare PTA observations with theoretical models. We find that GWB spectral inferences used by PTAs could be biased toward shallower slopes and higher amplitudes at $f=1/\rm yr$, thereby reducing the apparent tension between astrophysical expectations and PTA observations.
@@ 1 Introduction
Recently, various PTA experiments have presented the first observational evidence for a GW signal (NANOGrav, Agazie et al. 2023a; EPTA & InPTA, EPTA Collaboration et al. 2023b; PPTA, Reardon et al. 2023; CPTA, Xu et al. 2023; and MeerKAT, Miles et al. 2024). This signal shows both spatial and temporal correlations among pulsars, consistent with those expected from a stochastic GW background (Hellings and Downs, 1983). The low signal-to-noise ratio (S/N) of the detection makes the interpretation difficult, but one possibility is the incoherent superposition of GWs from a population of massive black hole binaries (MBHBs) with separations on the milliparsec scale. Indeed, inference of the GW background spectrum using PTA data confirms that the observed signal is broadly consistent with expectations from the MBHB population (Agazie et al., 2023c; EPTA Collaboration et al., 2024a), although the data tend to favor a larger amplitude than what has been theoretically predicted. One possibility is that the signal emanates from individually resolvable MBHBs that are particularly close and heavy, therefore standing above the background signal (e.g., Mingarelli et al., 2017; Agazie et al., 2023b). However, so far, this scenario is not supported by current PTA datasets (EPTA Collaboration et al., 2024b; Agazie et al., 2023b). Meanwhile, unmodeled noise or overly agnostic priors for the noise parameters could also explain part of the tension (Zic et al., 2022; Goncharov and others, 2025; van Haasteren, 2025). In this work, we show that an overly direct comparison between the theoretical power spectrum predicted by cosmological simulations (Kelley et al., 2017b; Sykes et al., 2022; Chen et al., 2025) or semi-analytical models (Wyithe and Loeb, 2003; McWilliams et al., 2014; Bonetti et al., 2018) and the data-inferred power spectra could lead to misleading conclusions about potential tensions between the two. In particular, the impact of the finite duration of the datasets on spectral inference should be considered more carefully [^1] After the submission of this paper, Crisostomi et al. (2025) also investigated the impact of improperly accounting for data windowing in PTA spectral inference..
@@
In the absence of a stronger GW detection, complementary data are necessary to understand the origin of the signal. Cosmological simulations can be used to understand the prevalence and nature of MBHBs in the Universe. These would also help in identifying which of these would be detectable with the PTAs and contribute to the GW background and/or be detectable as individual objects. Electromagnetic observations of these objects would help confirm the nature of the GW sources, but simulations can help us better understand where to search and provide the best strategy for the wavelength ranges to target (see e.g. Cella et al., 2025). With such prior knowledge, the overall challenging multi-messenger detection of MBHBs using PTAs (Charisi et al., 2022; Petrov et al., 2024) could be facilitated (Liu and Vigeland, 2021; Truant et al., 2025).
@@
In this paper, we use the Horizon-AGN cosmological simulation to identify and study the MBHBs that contribute most to the GW signal in the PTA band, either by contributing to the background signal or as resolvable continuous wave sources. We study the properties of these MBHBs along with those of their host galaxies. We determine the expected electromagnetic emission from radio to X-ray wavelengths, which we validate by comparing with the properties of MBHB candidates. We also identify the types of searches that would have the best chance of detecting the electromagnetic counterparts of the PTA sources. Finally, we investigate the commonly used methodology for comparing PTA observations with the predicted characteristic strain signal, highlighting its potential limitations when used to interpret the origin of the observed signal.
@@
The paper is organized as follows. Section 2 presents the Horizon-AGN simulation used in our study. In Sect. 3, we describe how we built our catalog of merging MBHBs from the simulation and modeled their sub-parsec dynamics in post-processing. In Sect. 4, we present how we generated realizations of the inspiralling MBHB population and inferred the associated characteristic strain signal. We also present the properties of the most contributing binaries and estimate their number in the PTA band. In Sect. 5, we compare the standard methodology used to convert the GW strain signal into timing residuals measured by PTAs, with a more accurate method, and highlight the potential limitations of the former. Using two toy PTA sensitivities, we also identified potentially resolvable individual binaries, for which we analyzed their properties and those of their host galaxies in Sect. 6, along with their electromagnetic properties, and evaluate their potential detectability in Sect. 7. We summarize our findings in Sect. 8.
@@ 2 The Horizon-AGN simulation
The gas in Horizon-AGN has an equation of state for an ideal monoatomic gas with an adiabatic index of $\gamma_{\rm ad}=5/3$. Gas cooling can be modeled down to a floor temperature of $10^{4}\,\rm K$ using the curves from Sutherland and Dopita (1993). Heating of the gas from a uniform UV background starts after redshift $z_{\rm reion}=10,$ following the prescription in Haardt and Madau (1996). In regions that exceed a gas hydrogen number density threshold of $n_{0}=0.1\,\rm H\,cm^{-3}$, star formation is triggered in a Poisson random process (Rasera and Teyssier, 2006; Dubois and Teyssier, 2008), following the Schmidt relation with a constant star formation efficiency of $\varepsilon_{*}=0.02$ (Kennicutt, 1998; Krumholz and Tan, 2007). A mechanical energy injection from Type Ia supernovae (Type Ia SNe), Type II SNe, and stellar winds is included assuming a Salpeter (1955) initial mass function (IMF) with cutoffs at $0.1\,{\rm M_{\odot}}$ and $100\,{\rm M_{\odot}}$.
@@
Massive black holes and their feedback follow the numerical implementation of Dubois et al. (2012). They are seeded in cells where the gas density is larger than $n_{0}$ and where the gas velocity dispersion is larger than $100\,\rm km\,s^{-1}$. The MBH seeding is stopped at $z=1.5$. To avoid formation of multiple MBHs in the same galaxy, an exclusion radius of 50 comoving kpc is imposed in the seeding process. To compensate for the inability to capture the multiphase nature of the interstellar gas, the MBH accretion rate is set to be the Bondi-Hoyle-Littleton rate multiplied by a factor $\alpha_{\rm b}=(n/n_{0})^{2}$ when $n>n_{0}$ and $\alpha_{\rm b}=1$ otherwise (Booth and Schaye, 2009). The effective accretion rate onto MBHs is capped at the Eddington luminosity with a radiative efficiency of 0.1. 15% of the MBH emitted energy is isotropically coupled to the gas within $4\Delta x$ as thermal energy for luminosities above 1% of the Eddington luminosity. If the 1% of the Eddington luminosity threshold is not reached, then the feedback takes a mechanical form instead, with 100% of the power injected into a bipolar cylindrical jet with a radius of $\Delta x$ and a height of $2\Delta x$, at a velocity of $10^{4}\,\rm km\,s^{-1}$.
@@
Instead of constantly repositioning MBHs at the minimum of the local potential, which causes unnatural dynamics (Tremmel et al., 2015), gas dynamical friction is exerted on the MBH to avoid spurious motions due to finite force resolution effects (see Dubois et al., 2013 for additional details). The boost factor used here is the same as the boost factor $\alpha_{\rm b}$ for accretion. The equation of gas dynamical friction is $F_{\rm DF}=f_{\rm gas}4\pi\alpha_{\rm b}\rho_{\rm gas}(GM_{\rm BH}/\bar{c}_{s})^{2}$, where $G$ is the gravitational constant, $\rho_{\rm gas}$ is the mass-weighted mean gas density within a sphere with a radius of $4\,\Delta x$, and $f_{\rm gas}$ is a factor function of the Mach number, ${\mathcal{M}}=\bar{u}/\bar{c}_{s}$, which accounts for the extension and shape of the wake (Ostriker, 1999). Finally, $f_{\rm gas}$ is in a range between 0 and 2 for an assumed Coulomb logarithm of 3 (Chapon et al., 2013; Lescaudron et al., 2023).
@@
In Horizon-AGN the dark matter halos and galaxies are identified with the AdaptaHOP halo finder (Aubert et al., 2004). The density field used in AdaptaHOP is smoothed over 20 particles. The identification density threshold is set to be 178 times the average total matter density. On top of the density criteria, either 50 dark matter particles or 50 star particles are required for a dark matter halo or galaxy identification. The centers of halos and galaxies are estimated using the shrinking sphere approach proposed by Power et al. (2003).
@@ 3 Selection and modeling of MBH binaries in Horizon-AGN
We generated a catalog of MBH mergers using the post-processed simulation data. To find merging MBHs, we searched through the data of all MBHs at each coarse time step ($\sim 0.6-0.7$ Myr) of the Horizon-AGN simulation. We note that this time step is much shorter than the time step over which a full output is saved, which is $\sim 150$ Myr. Each MBH was assigned a unique ID number when seeded. This ID is carried by the MBH through the simulation until it numerically merges with another MBH at a separation of $4\Delta x$. The merged MBH inherits the ID of the primary black hole in the merger, while the secondary MBH ID is erased after the merger. By searching for disappearing MBH IDs, we can find the corresponding MBH mergers. Possible spurious numerical mergers are filtered out by selecting only MBH pairs within $\max(2R_{\rm eff},4\Delta x)$ from the galaxy centers, where $R_{\rm eff}$ is the effective radius of the galaxy measured as its projected two-dimensional (2D) radius containing half of the galaxy stellar mass. This check was performed for the primary MBH at the outputs before the merger and for the merged MBH at the output after the merger. The reason is that the MBHs are likely to move relative to the galaxy center in the $150$ Myr between outputs.
@@
First, we added a dynamical friction phase from the position of the MBHs when they are numerically merged down to the point when they become gravitationally bound. The dynamical friction timescale is estimated assuming the MBH is in an isothermal sphere, considering only the stellar component of the galaxy and including a factor of $0.3$ to account for typical orbits being noncircular,
@@
where $M_{\rm BH}$ is the black hole mass, $\sigma_{\star}$ is the central stellar velocity dispersion approximated as $(0.25GM_{\star}/R_{\rm eff})^{1/2}$, $\Lambda=\ln(1+M_{\star}/M_{\rm BH})$, with $M_{\star}$ being the total stellar mass of the galaxy hosting the MBH at the output before the numerical merger and $d$ its distance from the galactic center. As confirmed by Li et al. (2020a), Li et al. (2020b), and Chen et al. (2022), the stellar dynamical friction dominates over the dynamical friction from gas or that from dark matter. We calculated the dynamical friction delay time of both MBHs in a pair and used the longer, which is normally associated with $M_{2}$.
We also accounted for the shrinking of the binary orbit until coalescence via stellar hardening, viscous torques in a circumbinary disc, and emission of gravitational waves, following Sesana and Khan (2015) and Dotti et al. (2015). The binary evolution timescale to coalescence was taken to be the minimum between the two following equations:
@@
Here, $q=M_{2}/M_{1}$ is the mass ratio of the MBHB. In Eq. 2, $\sigma_{\rm inf}$ and $\rho_{\rm inf}$ are the velocity dispersion and stellar density at the sphere of influence, defined as the sphere containing twice the binary mass in stars:
@@
where we continue to assume a singular isothermal sphere power-law density profile with slope 2-2, and $a_{\rm gw}$ is the separation at which the binary spends most of the time (see Sesana and Khan, 2015):
@@
The spatial resolution of the Horizon-AGN simulation does not provide reliable information on the density profile at such small orbital separations. However, it was demonstrated in (Volonteri et al., 2020) that the choice of a single isothermal sphere is in very good agreement with the densities measured in observations of local galaxies. In Eq. 3, $\varepsilon_{0.1}$ is the radiative efficiency normalized to 0.1. We followed Dotti et al. (2015) in selecting $a_{\rm i}=GM_{12}/2\sigma_{\star}^{2}$ and $a_{\rm c}=1.9\times 10^{-3}(M_{12}/10^{8}\,{\rm M_{\odot}})^{3/4}\,{\rm pc}$.
To model the eccentricity evolution in post-processing, we first assigned an eccentricity to each binary when reaching the influence radius, $r_{\rm inf}$, according to the eccentricity distribution shown in the right panel of Fig. 6 in Li et al. (2020a). The eccentricity is then evolved under loss-cone scattering, viscous drag from the circumbinary disk, and GW emission, from the influence radius to the final coalescence. The orbit hardening and eccentricity evolution due to the loss-cone scattering is described by
@@
where $f_{\rm orb}$ is the orbital frequency, and $H$ and $K$ are numerical factors resulting from the three-body scattering experiments (Quinlan, 1996; Sesana et al., 2006).
The viscous drag due to the circumbinary disk is included when the separation between the binary is below $1\,\rm pc$. Haiman et al. (2009) described how the binary orbit embedded in a circumbinary (Shakura and Sunyaev, 1973) $\alpha$-disk evolves due to viscous drag and how this evolution depends on the different physical conditions within the disk.
According to Haiman et al. (2009), there are different regimes for an MBHB in a gap-opened, $\alpha$-disk depending on: (1) whether the radiation pressure or gas pressure balance the vertical gravity ($r^{\rm gas/rad}$); (2) whether the opacity is dominated by electron scattering or free-free absorption ($r^{\rm es/ff}$); (3) whether the binary is massive enough compared to the local disk mass ($M_{\rm 2}$-dominated or disk-dominated). The characteristic radii are defined as
@@
According to Shapiro and Teukolsky (1983), a disk can be divided into three regions: (i) inner region ($r<r^{\rm gas/rad}$) with radiation-dominated pressure and electron scattering-dominated opacity; (ii) middle region ($r^{\rm es/ff}>r>r^{\rm gas/rad}$) with gas-dominated pressure and electron scattering-dominated opacity; and (iii) outer region ($r>r^{\rm es/ff}$) with gas-dominated pressure and free-free scattering-dominated opacity. Within each of those three region, there are two possibilities: 1) an $M_{\rm 2}$-dominated region ($r<r^{\rm\nu/s}$) and 2) a disk-dominated region ($r>r^{\rm\nu/s}$). The $r^{\rm\nu/s}$ in three regions of an $\alpha$-disk are defined in Haiman et al. (2009) as
@@
Thus, there are six regimes and their orbital frequency evolution rates are listed below.
@@
if $r^{\rm gas/rad}<r<r^{\rm es/ff}$ and $r>r_{\rm mid}^{\rm\nu/s}$; where $r_{\rm 3}$ is the orbital semi-major axis in units of $10^{\rm 3}R_{\rm sch}$, $q_{\rm s}=4q/(1+q)^{\rm 2}$ is the symmetric mass ratio. We note that this prescription implies that the MBHB orbit always shrinks under the influence of viscous drag, especially in the presence of stellar hardening suggested by some most recent simulations (Cuadra et al., 2009; Roedig et al., 2012; Bortolas et al., 2021; Franchini et al., 2021; Amaro-Seoane et al., 2023)